A computer literacy survey conducted by the ministry of health among a random sample of 1000 skilled employees in Kenya showed that 40% of them are computer literate. Based on this information, determine the 95% confidence interval for computer literacy among employees in Kenya.
1
Expert's answer
2020-09-20T18:00:08-0400

The formula for confidence interval is given by,


"p\ \pm \ Z\alpha*\sqrt{\frac{p(1-p)}{n}}" .............(Eq. 1)


where,

p = probability of employees who are computer literate from random selection of samples

n = sample size


In our case,

p = 40% = 0.4

n = 1000

"\alpha" = 1-95%= 5%


Since it is a 2 tailed test the value of Z = Z"\alpha"/2=Z0.025


From the standard normal distribution table we get the value of Z0.025 = 1.96


Substituting the values in equation 1 we get,


"0.4\ \pm\ 1.96\sqrt{\frac {0.4(1-0.4)}{1000}} = 0.4\pm0.0155"


Hence the confidence interval for literacy rate is (0.3845, 0.4155)